# Extracting Membrane-like hexagonal Boron Nitride hosting single Defect Centers for Resonator Integration

### v1

@@ Abstract
The integration of membranes into optical resonators plays a key role in a variety of applications, including optomechanics. If such membranes host atom-like systems, ideally with access to spin states, new roads in quantum photonics and also in optomechanics can be taken. Layered, two-dimensional materials have emerged as candidates for membranes hosting atom-like quantum emitters. Hexagonal boron nitride (hBN) is among the most promising two-dimensional platforms showing good mechanical properties combined with the ability to host various kinds of optical active (spin-) defects. However, the determinisitc creation of optically active defect centers in hBN membranes is an outstanding challenge. Commercially available flakes of hBN host defect centers with promising optical properties, but the integration into optical resonators suffers from scattering losses due to the flakes topography and suitable transfer, handling and manipulation techniques need to be established. Here, we develope a toolset of nano-scaled manipulation techniques to extract membrane-like structures of commercially-available hBN containing spectrally narrow single photon emitters. We demonstrate the transfer and integration into photonic devices, by coupling a single photon emitter in membran-like hBN to the mode of an open Fabry-Perot fiber cavity (FPFC) and observe cavity induced spectral enhancement by a factor of up to 100 at room temperature. Overcoming hBN-induced scattering for extracted hBN membranes, which host single photon emitters, paves the way for future applications such as its use as an optomechanical system.
@@ 1 Introduction
The integration of optomechanical membranes, based on silicone nitride (SiN) or graphene, into high-finesse Fabry Perot cavities has led to advancements in the field of optomechanics 1, 2, 3, 4, 5, 6. Membranes have a characteristic thickness signifcantly smaller than the optical wavelength. The much larger lateral dimensions can be tailored towards specific requirements such as
@@
emission rate and a reduced linewidth, in this case up to a factor of 100 and 160 respectively. These effects are attributed to cavity funneling which is enabled by the operation of the coupled system in the absence of scattering losses at the coating-limited finesse . All these metrica are useful for developments of future optomechanical experiments with access to atomic transitions or spin states.
@@ 2 Extraction of membrane-like hBN structures
As a first step a source sample of hBN flakes is prepared, where a commercially available emulsion of hBN (2D Semiconductors) is spin-coated onto a fused silica substrate after treatment in an ultrasonic bath. The spin-coated sample is annealed at 800 °C for 30 min under vacuum to improve the optical properties of optical active emitters 55. As a next step the sample is investigated in a home-built confocal microscope to pre-select emitters, suitable for further integration into a FPFC according to their spectral properties. An exemplary emitter spectrum is represented in fig. 1a) (grey). To investigate the topographic properties of the pre-selected emitters host, we conduct AFM scans.
As presented exemplary in fig. 1b) (grey) these scans reveal the topography of the host material, which is in this case an ensemble of a multitude of individual flakes. It is speculated that only one flake is hosting the emitter (excluding the possibility of the emitter being hosted between two or more particles). We therefore dismantle ensembles into membrane-like sub-ensembles and separate them by a few micrometers (fig. 1 (blue, orange and green)) with an AFM by driving the cantilever into contact with the individual flakes. The sub-ensembles are investigated again for their optical properties (as shown exemplary for membrane-like structures A and B in fig. 1 c)). A membrane-like character, with a thickness below nm (A) and 100 nm (B), thereby keeping the lateral extension in the micrometer range. In case of A, the thickness of the membrane-like structure is reduced to below a tenth of the emission wavelength of the emitter, making it particularly interesting for the integration into scattering sensitive micro-cavities. The membrane-like structures hosting emitters can then be transferred to a target sample as described in the next section. Particles without optical-active defect centers can be discarded (e.g. membrane-like structure C in fig. 1 b)) and remain on the source sample.
@@ 3 Membrane Transfer
After suitable defect centers in hBN are selected, they can be integrated into optical and photonic devices (here onto a macroscopic mirror of a FPFC). The pick- and place transfer technique leverages a home-built nano-manipulation system which consists of a tungsten-tip (tip radius < 0.1 µm) driven by a 3 axis piezo stage (fig. 2a),b) ). An additional home-built observation (wide-field) microscope on top of the sample provides optical feedback over the tip position (as illustrated in fig. 2a)). The combination with a confocal microscope enables at the same time near real-time optical characterization of individual flakes and membranes. The position of the tip can be monitored by both microscopes respectively. The tungsten-tip allows to lift off individual clusters of hBN flakes as well as membrane-like structures from the sample and transfer them to a target substrate, here a macroscopic cavity mirror (schematically illustrated in fig. 2 d)-e)). Focused ion beam (FIB) engraved marker patterns on the target mirror enable the assignability of individual emitters after a multitude of transfer processes. The compatibility of this technique with other target substrates enables the integration of hBN nanoparticles containing single photon emitters in a wide range of photonic and microscopic scaled applications 54. Subsequent confocal microscope investigation on the target substrate ensures the continued existence of the emitter(s). With the technique described, we transfer a multitude of flakes and membranes (>30) with different spectral and topographic properties onto a macroscopic cavity mirror and investigate them with a home-built scanning cavity microscope.
Similar to the separation method described in section 2, membrane-like structures of hBN can be extracted by driving the tungsten-tip into contact and applying mechanical shifts to the ensemble. The lack of resolution compared to the AFM technique is compensated with the real-time feedback of the confocal and observation microcope. Subsequent AFM scans can then reveal the topographic features of the extracted particles if required. For visualization purpose, we de-cluster an exemplary ensemble of flakes in a scanning electron microscope (SEM) with the same tungsten-tip as used the home-built optical setup as described in the supplementary material section 1.
@@ 4 Cavity Integration
Our target system is an open FPFC which consists of a microscopic curved mirror on the tip of an optical single mode fiber and a macroscopic planar mirror. The conical mirror on the tip of the fiber is machined by a combination of FIB milling and CO2 laser smoothing (ROC µm) 56 and coated with a dielectric coating at Laseroptik GmbH yielding a coating-limited finesse (between nm and nm). The planar mirror is equipped with multiple membrane-like structures and flakes of hBN containing defect centers as described in the previous sections. The cavity is enclosed in a vacuum chamber combined with a home-built vibration isolation stage to minimize the impact of acoustic and thermal noise onto the system. Piezo driven slick stick actuators (Attocube Systems AG) allow the precise positioning of the macroscopic planar mirror relative to the cavity mode. Scanning the planar mirror in the lateral direction of the mirror enables the system to act as a scanning cavity microscope 57. When illuminated in transmission via the fiber port, FIB engraved marker patterns as well as transferred hBN flakes and membrane-like structures can be resolved as presented in fig. 3b). Darker regions correlate to higher optical losses (mostly caused by scattering) compared to brighter regions, which is why membrane-like structures with low scattering might be not resolvable with this technique. AFM scans of flake D and membrane-like structure E confirm that an increased particle size correlates to an increased extinction rate. The indication of low optical losses induced by scattering or absorption as well as the narrow emission line (nm, fig. 3c)) makes emitter E a promising candidate for coupling to the optical mode of the cavity. Once an interesting particle or membrane is selected, a piezo allows the control of the cavity length , enabling active stabilization of the resonator with a Pound-Drever-Hall locking scheme (as schematically illustrated in fig. 3 a)).
@@ 5 Coupled System
Coupling the optical dipole of the defect center in a transferred membrane-like structure (fig. 3 E) to the optical mode of the FPFC results in a narrowed and enhanced emission compared to the free-space emission. The relative position between the cavity fiber and the planar mirror are tuned with a piezo until a cavity mode and the emitter’s ZPL are spectrally overlapped (fig. 4 a)). Since the emission linewidth of the coupled system is below the resolution limit of our spectrometer, the linewidth of the cavity system on the emitter is determined by locking the cavity with a laser at the next higher TEM00 cavity resonance. The resonance on the emitter is probed with a second laser (Dye ring laser system) as presented in fig. 4 b). After fitting a Lorentzian function to the transmission data, a locked linewidth of 8.9 GHz is obtained, which corresponds to a spectral narrowing factor of 160 compared to the free space emission of the emitter. Tuning the length of the cavity allows us to probe and reconstruct the free space emission spectrum of the emitter (depicted normalized in blue in fig. 4 a)). We compare the absolute free space emission of the emitter with the cavity coupled emission using the same optical setup and components, only flipping the mirror surface towards the objective. The spectral density can be calculated for both cases (free space emission and cavity coupled emission) when exciting the system off-resonantly at nm, which yields a 100-fold enhanced spectral density (fig. 4 c)) in cavity configuration. This enhancement is caused by cavity funneling of the thermally broadened emission (bad emitter regime) into the cavity mode 58, 59, 60. Second-order auto correlation measurements (fig. 4 d)) reveal the single-photon emitting character of the defect center after fitting an exponential function with . Pulsed lifetime measurements indicate an optical lifetime of the coupled system of ns which is comparable to the free space lifetime (as seen in the supplementary material section 2.). Due to the operation at room temperature the emission is thermally broadened resulting in a reduced quality factor 51. This quality factor is lower compared to the cavity quality factor (bad emitter regime).
The resulting effective Purcell factor is estimated by
@@
where denotes the measured finesse of the emitter-cavity system and denotes the finesse of the empty cavity. Both are depicted in fig. 4g) for different wavelengths together with the coating limited finesse given by the individual transmission of the cavity mirrors (measured by Laseroptik Gmbh).

### v2

@@ Abstract
The integration of membranes into optical resonators plays a key role in a variety of applications, including optomechanics. If such membranes host atom-like systems, ideally with access to spin states, new roads in quantum photonics and also in optomechanics can be taken.
Layered, two-dimensional materials have emerged as candidates for membranes hosting atom-like quantum emitters. Hexagonal boron nitride (hBN) is among the most promising two-dimensional platforms showing good mechanical properties combined with the ability to host various kinds of optical active (spin-) defects. However, the determinisitc creation of optically active defect centers in hBN membranes is an outstanding challenge. Commercially available flakes of hBN host defect centers with promising optical properties, but the integration into optical resonators suffers from scattering losses due to the flakes topography and suitable transfer, handling and manipulation techniques need to be established. Here, we develope a toolset of nano-scaled manipulation techniques to extract membrane-like structures of commercially-available hBN containing spectrally narrow single photon emitters. We demonstrate the transfer and integration into photonic devices, by coupling a single photon emitter in membran-like hBN to the mode of an open Fabry-Perot fiber cavity (FPFC) and observe cavity induced spectral enhancement by a factor of up to 100 at room temperature. Overcoming hBN-induced scattering for extracted hBN membranes, which host single photon emitters, paves the way for future applications such as its use as an optomechanical system.
Keywords: 2D-materials, Fabry-Perot fiber cavities, hexagonal boron nitride, solid state defect centers
@@ 1 Introduction
The integration of optomechanical membranes, based on silicone nitride (SiN) or graphene, into high-finesse Fabry-Perot cavities has led to advancements in the field of optomechanics 1, 2, 3, 4, 5, 6. Membranes have a characteristic thickness signifcantly smaller than the optical wavelength. The much larger lateral dimensions can be tailored towards specific requirements such as
@@
emission rate and a reduced linewidth, in this case up to a factor of 100 and 160 respectively. These effects are attributed to cavity funneling which is enabled by the operation of the coupled system in the absence of scattering losses at the coating-defined finesse . All these metrica are useful for developments of future optomechanical experiments with access to atomic transitions or spin states.
@@ 2 Extraction of membrane-like hBN structures
As a first step a source sample of hBN flakes is prepared, where a commercially available emulsion of hBN (2D Semiconductors) is spin-coated onto a fused silica substrate after treatment in an ultrasonic bath. The spin-coated sample is annealed at 800 °C for 30 min under vacuum to improve the optical properties of optical active emitters 55. As a next step the sample is investigated in a home-built confocal microscope to pre-select emitters, suitable for further integration into a FPFC according to their spectral properties. An exemplary emitter spectrum is represented in fig. 1a) (grey). To investigate the topographic properties of the pre-selected emitters host, we conduct AFM scans. As presented exemplary in fig. 1b) (grey) these scans reveal the topography of the host material, which is in this case an ensemble of a multitude of individual flakes. It is speculated that only one flake is hosting the emitter (excluding the possibility of the emitter being hosted between two or more particles). We therefore dismantle ensembles into membrane-like sub-ensembles and separate them by a few micrometers (fig. 1 (blue, orange and green)) with an AFM by driving the cantilever into contact with the individual flakes. The sub-ensembles are investigated again for their optical properties (as shown exemplary for membrane-like structures A and B in fig. 1 c)). A membrane-like character, with a thickness below nm (A) and 100 nm (B), thereby keeping the lateral extension in the micrometer range. In case of A, the thickness of the membrane-like structure is reduced to below a tenth of the emission wavelength of the emitter, making it particularly interesting for the integration into scattering sensitive micro-cavities. The membrane-like structures hosting emitters can then be transferred to a target sample as described in the next section. Particles without optical-active defect centers can be discarded (e.g. membrane-like structure C in fig. 1 b)) and remain on the source sample.
@@ 3 Membrane Transfer
After suitable defect centers in hBN are selected, they can be integrated into optical and photonic devices (here onto a macroscopic mirror of a FPFC). The pick- and place transfer technique leverages a home-built nano-manipulation system which consists of a tungsten-tip (tip radius < 0.1 µm) driven by a 3 axis piezo stage (fig. 2a),b) ). An additional home-built observation (wide-field) microscope on top of the sample provides optical feedback over the tip position (as illustrated in fig. 2a)).
The combination with a confocal microscope enables at the same time near real-time optical characterization of individual flakes and membranes. The position of the tip can be monitored by both microscopes respectively. The tungsten-tip allows to lift off individual clusters of hBN flakes as well as membrane-like structures from the sample and transfer them to a target substrate, here a macroscopic cavity mirror (schematically illustrated in fig. 2 d)-e)). Focused ion beam (FIB) engraved marker patterns on the target mirror enable the assignability of individual emitters after a multitude of transfer processes. The compatibility of this technique with other target substrates enables the integration of hBN nanoparticles containing single photon emitters in a wide range of photonic and microscopic scaled applications 54. Subsequent confocal microscope investigation on the target substrate ensures the continued existence of the emitter(s). With the technique described, we transfer a multitude of flakes and membranes (>30) with different spectral and topographic properties onto a macroscopic cavity mirror and investigate them with a home-built scanning cavity microscope. Similar to the separation method described in the previous section, membrane-like structures of hBN can be extracted by driving the tungsten-tip into contact and applying mechanical shifts to the ensemble. The lack of resolution compared to the AFM technique is compensated with the real-time feedback of the confocal and observation microcope. Subsequent AFM scans can then reveal the topographic features of the extracted particles if required. For visualization purpose, we de-cluster an exemplary ensemble of flakes in a scanning electron microscope (SEM) with the same tungsten-tip as used the home-built optical setup as described in the supplementary material section 1.
@@ 4 Cavity Integration
Our target system is an open FPFC which consists of a microscopic curved mirror on the tip of an optical single mode fiber and a macroscopic planar mirror. The conical mirror on the tip of the fiber is machined by a combination of FIB milling and CO2 laser smoothing (ROC µm) 56 and coated with a dielectric coating at Laseroptik GmbH yielding a target finesse of (between nm and nm). The planar mirror is equipped with multiple membrane-like structures and flakes of hBN containing defect centers as described in the previous sections. The cavity is enclosed in a vacuum chamber combined with a home-built vibration isolation stage to minimize the impact of acoustic and thermal noise onto the system.
Piezo driven slick stick actuators (Attocube Systems AG) allow the precise positioning of the macroscopic planar mirror relative to the cavity mode. Scanning the planar mirror in the lateral direction of the mirror enables the system to act as a scanning cavity microscope 57. When illuminated in transmission via the fiber port, FIB engraved marker patterns as well as transferred hBN flakes and membrane-like structures can be resolved as presented in fig. 3b). Darker regions correlate to higher optical losses (mostly caused by scattering) compared to brighter regions, which is why membrane-like structures with low scattering might be not resolvable with this technique. AFM scans of flake D and membrane-like structure E confirm that an increased particle size correlates to an increased extinction rate. The indication of low optical losses induced by scattering or absorption as well as the narrow emission line (nm, fig. 3c)) makes emitter E a promising candidate for coupling to the optical mode of the cavity. Once an interesting particle or membrane is selected, a piezo allows the control of the cavity length , enabling active stabilization of the resonator with a Pound-Drever-Hall locking scheme (as schematically illustrated in fig. 3 a)).
@@ 5 Coupled System
Coupling the optical dipole of the defect center in a transferred membrane-like structure (fig. 3 E) to the optical mode of the FPFC results in a narrowed and enhanced emission compared to the free-space emission. The relative position between the cavity fiber and the planar mirror are tuned with a piezo until a cavity mode and the emitter’s ZPL are spectrally overlapped (fig. 4 a)). Since the emission linewidth of the coupled system is below the resolution limit of our spectrometer, the linewidth of the cavity system on the emitter is determined by locking the cavity with a laser at the next higher TEM00 cavity resonance. The resonance on the emitter is probed with a second laser (Dye ring laser system) as presented in fig. 4 b). After fitting a Lorentzian function to the transmission data, a locked linewidth of 8.9 GHz is obtained, which corresponds to a spectral narrowing factor of 160 compared to the free space emission of the emitter.
Tuning the length of the cavity allows us to probe and reconstruct the free space emission spectrum of the emitter (depicted normalized in blue in fig. 4 a)). We compare the absolute free space emission of the emitter with the cavity coupled emission using the same optical setup and components, only flipping the mirror surface towards the objective. The spectral density can be calculated for both cases (free space emission and cavity coupled emission) when exciting the system off-resonantly at nm, which yields a 100-fold enhanced spectral density (fig. 4 c)) in cavity configuration. This enhancement is caused by cavity funneling of the thermally broadened emission (bad emitter regime) into the cavity mode 58, 59, 60. Second-order auto correlation measurements (fig. 4 d)) reveal the single-photon emitting character of the defect center after fitting an exponential function with . Pulsed lifetime measurements indicate an optical lifetime of the coupled system of ns which is comparable to the free space lifetime (as seen in the supplementary material section 2.). Due to the operation at room temperature the emission is thermally broadened resulting in a reduced quality factor 51. This quality factor is lower compared to the cavity quality factor (bad emitter regime). The resulting effective Purcell factor is estimated by
@@
where denotes the measured finesse of the emitter-cavity system and denotes the finesse of the empty cavity. Both are depicted in fig. 4g) for different wavelengths together with the coating-designed finesse given by the individual transmission of the cavity mirrors (measured by Laseroptik Gmbh).